OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y + 9)2 = 4. What is the center and radius of circle S? Center: (16, -9); Radius: 4 Center: (-16, 9); Radius: 2 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 4

OpenStudy (anonymous):

\[Is \it : (x-16)^2+(y+9)^2=4 \]

OpenStudy (anonymous):

The standard form of circle is \[(x-h)^2+(y-k)^2=r^2 \] where h and k are centre of circle and r is radius. now u can compare and will get ur answer